determine the maximum profit in the price that would yield the maximum profit for P=-400p^2+12,400p-50,000


1
Expert's answer
2021-09-22T17:53:56-0400

To find the price which would yield the maximum profit we divide b by 2a as follows:

P"=-" 400P"^{2}+" 12400P "-" 50000

"\frac{-b}{2a}=" "\frac{-12400}{2(-400)}"

"=\frac{-12400}{-800}"

Answer is 15.50

To find the maximum profit we substitute this value of price in the profit equation as follows;

P"=-" 400"{(15.5}^2)" "+" 12400"(" 15.5")" "-50000"

"-" 96100"+" 192200"-" 50000

Answer is 46 100

So the profit maximizing price is $15.50 and the maximum profit at this price is $46 100